From a worksheet:
A woman invests a total of $8550 in two accounts. One pays 8% in annual interest and the other pays 7.5% interest. If she made $666.63 in interest for the first year, how much is in each account?

Answer by checkley75(3666)

.08X+.075(8550-X)=666.63
.08X+641.25-.075X=666.63
.005X=666.63-641.25
.005X=25.38
X=25.38/.005
X=5076 INVESTED @ 8%
8550-5076=3474 INVESTED @ 7.5%
PROOF
.08*5076+.075*3474=666.63
406.08+260.55=666.63
666.63=666.63